OpenStudy (anonymous):

What is 2 to the power of 3 over 2 equal to? square root of 8 cube root of 8 cube root of 16 square root of 16

OpenStudy (anonymous):

@mathmale please help me (:

OpenStudy (anonymous):

\[2^{\frac{3}{2}}=\sqrt{2^3}=\sqrt{8}\]

OpenStudy (anonymous):

thank you so much ^-^

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oh wait im sorry the questions for some howq made it diff this is what is actually suppost to be 

OpenStudy (anonymous):

@JoannaBlackwelder please help me when u get the chance